OpenStudy (anonymous):

Explain how raising a quotient to a power is similar to raising a product to a power

OpenStudy (mertsj):

In both cases, you multiply the exponents.

OpenStudy (amistre64):

a quotient is a product is a quotient is a product ....

OpenStudy (amistre64):

\[\frac ab = ab^{-1}\] \[(\frac ab)^n = (ab^{-1})^n=a^nb^{-n}=\frac{a^n}{b^n} \]